Find all solutions of the congruences:

(e) 64x 83 (mod 105) (f) 589x 209 (mod 817) (g) 49x 5000 (mod 999)
Answer #1

e T given Here, (64, 10s)nd83Isy tee isexactly ee in conghuent selution modulo tos. 64--드 41x1 +23 41.= 23x1 tlg. 23 18 XI 5-1x23 + x (41-23) 3x41-162(64-y? CYL-3403) 2.- 33(mod-10s) 64 (-34D3)三83(mod 105) ウ . ok 62 (mool los)) Given conglu emte uenceA /9/209, [S89 ,814)こ19 rue ,8 tere aeomgkuen ce nconghuen New, เจ้ /9 31x-=-1L(mod 43) う Jn Euclide= 3X구-4 (12-7) 4x12+ 7 x (31-12x2)-- (93-31) = 7.x31-18 31x(25x11)三11 (mod 43) → It (mod 43) ond hence ot ea O , tto-19.. Aoltut Sch au はマ.ueAjlined. selations Given eongkuente i theu is exastly one in cogent selution me dulo 9 Poy. Euclioloan. Algor3--( 8-3(2) 3-8+3(2) 3(3)--8 3(11-3)-8 = = = ーーチ(49)-18(999-20[49)) E. t (49)-1 8 (999) t-860-1 41) (367)(99)土-(-18)(999) = 3三1 (mod 999) 1 (mod 999). зст x49 49(367X5000) Ξ S000 (Yn0d 999) ラー 49( 1935000)三5000 (mod 999) ウ os0c 1835000 med (199) med